OpenStudy (buggiethebug):

Determine the coordinates of two points of tangency to the curve y=-2x^2, given that the corresponding tangent lines intersect at point (2,8). Can someone please tell me what the answer is? and what the approach it took

OpenStudy (anonymous):

a point on the curve looks like \((x,-2x^2)\) and if the tangent line passes through \((2,8)\) it will have slope \[\frac{-2x^2-8}{x-2}\]

OpenStudy (anonymous):

it will also have the same slope as the derivative at that point, and the derivative is \(-4x\) meaning you have an equation \[\frac{-2x^2-8}{x-2}=-4x\] which you can solve for \(x\)

OpenStudy (anonymous):

probably easiest to start by multiplying both sides by \(-1\) and solving \[\frac{2x^2+8}{x-2}=4x\]

OpenStudy (anonymous):

pretty clear that you are going to get a quadratic equation to solve, which is why there will be two solutions

OpenStudy (buggiethebug):

wait can you tell me what the answers are, i have no clue, and there isn't a way to check

OpenStudy (buggiethebug):

we didnt learn derivatives...

OpenStudy (buggiethebug):

are the x values 2+/-2sqrt(2) ?

OpenStudy (anonymous):

you can't possibly do this problem without derivatives at least no way that i can think of

OpenStudy (anonymous):

yes you have the right answers, \[x=2\pm\sqrt{2}\]